# Federated Time Series Generation on Feature and Temporally Misaligned Data

### v2

@@ 20 Finetune $\mathcal{D}$ on updated $\mathbf{X}^{\text{pub}}$
[…]
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the balance between time and frequency losses while $w_{t}$ emphasizes learning at larger diffusion steps, with $\lambda$ being a small constant. The parameter $\delta_{t}\in(0,1)$ determines the amount of noise added at each forward diffusion step, where $t$ is a diffusion time step uniformly sampled from 111 to $T$ during training. The cumulative product $\bar{\gamma}_{t}=\prod_{v=1}^{t}\gamma_{v}$, with $\gamma_{t}=1-\delta_{t}$, track how the original signal diminishes over time due to the added noise. By weighting the loss at different steps, $w_{t}$ helps the model focus on reconstructing the signal under high-noise conditions.
@@
[…]
@@ 4.1 Time Series Generation
Context-FID Centralized* 0.682$\pm$0.106 0.281$\pm$0.040 0.782$\pm$0.138 0.533$\pm$0.082 1.737$\pm$0.125
Centralized 3.548$\pm$0.990 8.870$\pm$2.295 10.00$\pm$2.814 9.343$\pm$2.808 13.56$\pm$3.357
Local 1.648$\pm$0.229 1.313$\pm$0.188 0.751$\pm$0.121 1.179$\pm$0.179 1.694$\pm$0.153
Pre-trained 1.047$\pm$0.169 0.326$\pm$0.040 0.617$\pm$0.090 0.412$\pm$0.054 1.411$\pm$0.102
FedTDD 0.675$\pm$0.087 0.271$\pm$0.038 0.529$\pm$0.068 0.376$\pm$0.056 1.459$\pm$0.099
Correlational Centralized* 0.061$\pm$0.043 0.253$\pm$0.094 1.989$\pm$0.247 5.231$\pm$1.294 7.900$\pm$0.384
Centralized 0.769$\pm$0.336 0.340$\pm$0.097 2.230$\pm$0.518 5.681$\pm$0.634 18.07$\pm$2.311
Local 0.156$\pm$0.120 0.239$\pm$0.079 1.298$\pm$0.260 3.447$\pm$0.838 5.992$\pm$0.383
Pre-trained 0.077$\pm$0.052 0.165$\pm$0.074 1.323$\pm$0.171 2.821$\pm$0.651 6.049$\pm$0.349
FedTDD 0.058$\pm$0.050 0.161$\pm$0.064 1.296$\pm$0.215 2.800$\pm$0.686 6.017$\pm$0.364
Discriminative Centralized* 0.136$\pm$0.091 0.199$\pm$0.061 0.297$\pm$0.108 0.230$\pm$0.080 0.422$\pm$0.074
Centralized 0.476$\pm$0.042 0.475$\pm$0.017 0.474$\pm$0.024 0.496$\pm$0.006 0.477$\pm$0.030
Local 0.340$\pm$0.153 0.298$\pm$0.060 0.200$\pm$0.092 0.329$\pm$0.087 0.397$\pm$0.061
Pre-trained 0.175$\pm$0.117 0.115$\pm$0.060 0.208$\pm$0.068 0.141$\pm$0.068 0.419$\pm$0.051
FedTDD 0.185$\pm$0.105 0.106$\pm$0.061 0.153$\pm$0.120 0.153$\pm$0.072 0.414$\pm$0.051
Predictive Centralized* 0.040$\pm$0.000 0.127$\pm$0.003 0.112$\pm$0.015 0.292$\pm$0.009 0.137$\pm$0.004
Centralized 0.047$\pm$0.012 0.223$\pm$0.020 0.165$\pm$0.060 0.427$\pm$0.053 0.233$\pm$0.051
Local 0.043$\pm$0.003 0.118$\pm$0.011 0.048$\pm$0.006 0.204$\pm$0.012 0.135$\pm$0.006
Pre-trained 0.046$\pm$0.001 0.104$\pm$0.004 0.052$\pm$0.004 0.177$\pm$0.005 0.133$\pm$0.006
FedTDD 0.041$\pm$0.001 0.101$\pm$0.004 0.048$\pm$0.004 0.175$\pm$0.006 0.133$\pm$0.004
@@ 4.2 Ablation Study
Context-FID Centralized* 0.682$\pm$0.106 0.281$\pm$0.040 0.782$\pm$0.138 0.533$\pm$0.082 1.737$\pm$0.125
Centralized 3.733$\pm$0.959 11.54$\pm$3.894 14.68$\pm$4.263 13.17$\pm$3.035 15.34$\pm$4.789
Local 1.982$\pm$0.234 0.824$\pm$0.105 0.660$\pm$0.100 0.844$\pm$0.127 1.220$\pm$0.098
Pre-trained 0.738$\pm$0.142 0.316$\pm$0.032 0.547$\pm$0.099 0.381$\pm$0.066 1.178$\pm$0.104
FedTDD 0.680$\pm$0.123 0.267$\pm$0.036 0.510$\pm$0.072 0.331$\pm$0.051 1.196$\pm$0.098
Correlational Centralized* 0.061$\pm$0.043 0.253$\pm$0.094 1.989$\pm$0.247 5.231$\pm$1.294 7.900$\pm$0.384
Centralized 0.697$\pm$0.168 0.523$\pm$0.095 2.317$\pm$0.597 5.781$\pm$0.924 31.35$\pm$4.923
Local 0.091$\pm$0.052 0.167$\pm$0.057 1.079$\pm$0.196 1.984$\pm$0.594 4.929$\pm$0.395
Pre-trained 0.028$\pm$0.027 0.132$\pm$0.054 1.115$\pm$0.233 1.795$\pm$0.577 5.033$\pm$0.323
FedTDD 0.025$\pm$0.022 0.137$\pm$0.064 1.060$\pm$0.209 1.737$\pm$0.282 5.005$\pm$0.317
Discriminative Centralized* 0.136$\pm$0.091 0.199$\pm$0.061 0.297$\pm$0.108 0.230$\pm$0.080 0.422$\pm$0.074
Centralized 0.475$\pm$0.041 0.469$\pm$0.020 0.479$\pm$0.026 0.494$\pm$0.010 0.484$\pm$0.023
Local 0.300$\pm$0.116 0.208$\pm$0.070 0.190$\pm$0.088 0.241$\pm$0.071 0.398$\pm$0.058
Pre-trained 0.119$\pm$0.088 0.116$\pm$0.067 0.163$\pm$0.088 0.130$\pm$0.058 0.418$\pm$0.050
FedTDD 0.112$\pm$0.097 0.107$\pm$0.078 0.157$\pm$0.104 0.120$\pm$0.067 0.412$\pm$0.057
Predictive Centralized* 0.040$\pm$0.000 0.127$\pm$0.003 0.112$\pm$0.015 0.292$\pm$0.009 0.137$\pm$0.004
Centralized 0.168$\pm$0.025 0.196$\pm$0.027 0.198$\pm$0.049 0.314$\pm$0.052 0.223$\pm$0.029
Local 0.084$\pm$0.038 0.114$\pm$0.009 0.069$\pm$0.010 0.199$\pm$0.007 0.130$\pm$0.005
Pre-trained 0.028$\pm$0.007 0.108$\pm$0.004 0.063$\pm$0.007 0.190$\pm$0.005 0.132$\pm$0.005
FedTDD 0.028$\pm$0.005 0.107$\pm$0.005 0.062$\pm$0.006 0.186$\pm$0.004 0.130$\pm$0.005

### v3

@@ 20 Finetune $\mathcal{D}$ on updated $\mathbf{X}^{\text{pub}}$
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, $\mathbf{X}^{\text{pub}}_{j,k}$ is the $(j,k)$-th entry of $\mathbf{X}^{\text{pub}}$, $\tilde{\mathbf{X}}^{\text{pub}}_{j,k}$ is the denoised estimate from $\mathcal{D}$, and FFT denotes the Fast Fourier Transform [8], which is a mathematical operation that converts a finite-length time domain signal to its frequency domain representation. We take the following objective
@@
where $\lambda_{1}$ and $\lambda_{2}$ control the balance between time and frequency losses while $w_{t}$ emphasizes learning at larger diffusion steps, with $\lambda$ being a small constant. The parameter $\delta_{t}\in(0,1)$ determines the amount of noise added at each forward diffusion step, where $t$ is a diffusion time step uniformly sampled from $1$ to $T$ during training. The cumulative product $\bar{\gamma}_{t}=\prod_{v=1}^{t}\gamma_{v}$, with $\gamma_{t}=1-\delta_{t}$, track how the original signal diminishes over time due to the added noise. By weighting the loss at different steps, $w_{t}$ helps the model focus on reconstructing the signal under high-noise conditions.
@@
where $\mathbf{X}^{i}_{\text{train}_{j,k}}$ is the $(j,k)$-th entry of $\mathbf{X}^{i}_{\text{train}}$, $\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}$ is the denoised estimate from $\mathcal{U}$. After training, each client uses the trained imputer $\mathcal{U}^{i}$ to generate a synthetic dataset through unconditional synthesis, which includes both the common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ and the exclusive features $\hat{\mathbf{X}}_{\text{ex}}^{i}$. Starting from Gaussian noise, the imputer generates samples $\hat{\mathbf{X}}^{i}=\mathcal{U}^{i}(\mathbf{z}),\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, that capture the distribution of both common and exclusive features.
@@ 4.1 Time Series Generation
[…]
@@ 4.2 Ablation Study
[…]